rvc (rvc):

iif Nx^2-5xy+6y^2+x-3y=0 represents a pair of straight lines,then their point of intersection is =?

OpenStudy (goformit100):

omg @pooja195 mA'AM can Help you out.

rvc (rvc):

@divu.mkr @nirmalnema @rational @ikram002p

OpenStudy (anonymous):

partial derivate wrt x and wrt y and solve both you will get the point of intersection

rvc (rvc):

haan wohi partial derivate nahi aatha

OpenStudy (anonymous):

x ko constant le aur y ko derivate kar le and reverse

rvc (rvc):

haan wait

rvc (rvc):

what about A

OpenStudy (anonymous):

A kya hai?

rvc (rvc):

actually woh lamda hai

OpenStudy (anonymous):

lambda kaha?

rvc (rvc):

A ke jaga

rvc (rvc):

i mean N ke jaga :p

OpenStudy (anonymous):

2x-5y+1=0 -5x+12y-3=0

OpenStudy (anonymous):

partial derivatives

OpenStudy (anonymous):

ek variable rakh na

rvc (rvc):

got it thanks